OpenStudy (anonymous):

If f(x)=-3√ x-13 find ( f22 ). A. -9 B. -27 C. -6 D. -12

OpenStudy (brimblings):

A

OpenStudy (anonymous):

plug in 22 for x. you then get f(22)=-3*sqrt (22-13)=-3*sqrt9=-3*3=-9